# Exploration and Incentivized Participation in Randomized Trials

### v11

@@ Abstract
Keywords: exploration-exploitation tradeoff, incentive-compatibility, economic design, randomized controlled trials, clinical trials, multi-armed bandits
@@ 1 Introduction
Our scope. We initiate the study of incentivized participation in RCTs. Our objective is to incentivize participation while optimizing statistical performance. We frame it as a non-standard variant of exploration-exploitation tradeoff: the RCT always prefers to explore as uniformly as possible, whereas each patient/user prefers to exploit (choose an arm that seems best), and might not participate if the RCT is inconsistent with this preference. [^5] The standard tradeoff has an algorithm balance exploration vs exploitation to converge on the best arm. We henceforth refer to the potential RCT participants as agents.
@@
Our model. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentive constraints. An RCT (mechanism) sequentially interacts with self-interested agents like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^7] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. Formally, the mechanism optimizes the worst-case estimation error for these counterfactual estimates, denoted $\mathtt{ERR}$. If not for the incentive constraints for participation, the mechanism would simply choose arms independently and uniformly at random for each agent.
The role of $f$, the RCT’s metric, is as follows. Typically, $f(\omega)$ corresponds to the “societal value” of the outcome $\omega$. The RCT may also be interested in the empirical frequency of $\omega$, which is captured by setting $f={\bf 1}_{\left\{\,\omega\,\right\}}$, the indicator function that chooses this outcome. Our definition of $\mathtt{ERR}$ handles all metrics $f$ at once, and therefore applies to estimating societal value as well as outcome frequencies. In contrast, agents have subjective utilities over the outcomes that they wish to optimize, possibly different from the “societal value” or any other metric used by the RCT.
We address participation incentives by imposing Bayesian Individual Rationality (BIR) constraints, a standard notion in economic theory. Specifically, the mechanism commits to some randomized policy that assigns an arm to each agent, depending on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (as discussed above). BIR states that each agent must prefer participation to her “outside option,” in expectation over her Bayesian beliefs.
We consider three model variants. To isolate the core issues of incentivized participation, we first consider the fundamental case of homogeneous agents with the same beliefs and preferences. Then, the general case of heterogeneous agents endows each agent with a type. Then observable attributes such as medical history or user profile comprise one’s public type and determine one’s beliefs over outcomes. Moreover, one has unobservable preferences over the outcomes (private type) which need to be elicited in an incentive-compatible way. [^8] In practice, preference elicitation would happen during registration/opt-in process. For online platforms, private preferences and their elicitation make sense only if the RCT has explicit opt-in. The latter is uncommon, but may be needed if the RCT involves high-stakes interactions (such as buying a house) or is potentially controversial. The third model variant provides the mechanism with estimated type frequencies over the agent population. This final variant addresses situations in which some types are inherently (much) more “difficult” to handle than others. When such “difficult” types are rare, we leverage the type frequencies to mitigate their harmful influence.
@@
Agents’ beliefs. While we allow adversarial outcomes, the agents believe that the outcome-generating process is much more well-behaved. Specifically, the agents believe that one’s outcome distribution is determined by the assigned arm and one’s public type (if any), independent of the round. We call such beliefs time-invariant. We posit that the beliefs are shared by all agents and are known to the mechanism, as an appropriately defined Bayesian prior.
@@
This dichotomy between adversarial outcomes and simpler, time-invariant beliefs of the agents is essential to our model. On the one hand, our statistical guarantees are model-free once the data is collected, matching the standard for RCTs. On the other hand, time-invariant beliefs are standard in modeling sequential decision-makers (e.g., in incentivized exploration, see Related Work). While they need not be correct, they are reasonable as a model of reality and quite realistic as beliefs that patients or platform users might have, particularly so for clinical RCTs where they represent pre-existing medical knowledge. Beliefs whose support does not include the actual problem instance (a.k.a. misspecified beliefs) are not uncommon in economic theory, see Related Work.
@@
Crucially, the BIR constraint only needs to hold with respect to agents’ time-invariant beliefs. Put differently, it suffices to provide participation incentives for the hypothetical world in which the agents’ beliefs are correct. We note that a stronger version of participation incentives — robust against an adversarially chosen outcome sequence — is impossible to achieve (cf. Footnote 16).
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage (warm-up data) creates information asymmetry in the main state which drives participation incentives, whereas the main stage determines the counterfactual estimates. RCTs with a warm-up stage are very common in practice: e.g., phase-I clinical trials to calibrate the dosage and “pilot experiments” in online platforms.
@@
We obtain a near-optimal solution for each model variant described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward an idealized benchmark which focuses on a single round of our problem. Then, we design a mechanism whose estimation error $\mathtt{ERR}$ (in our actual, multi-round model) is upper-bounded in terms of this single-round benchmark. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. While the benchmark, the mechanism, and the lower bound are different for each model variant, many ideas carry over from one model variant to the next.
@@
Our mechanisms have a simple structure detailed below (which, however, suffices to connect them to the single-round benchmark). First, their “sampling distributions” — the distributions from which the arms are sampled — do not adapt to the data collected in the main stage. This is a desirable property: indeed, clinical RCTs in practice typically do not adapt to data, except possibly at a few pre-determined “review points”, [^9] This is because treating the patients and/or observing the outcomes may take a long time, and the doctors may find it difficult to coordinate across the patients. So, many/most patients may need to be treated in parallel. and neither do most RCTs in online platforms due to engineering considerations. Second, the type-dependent sampling distribution does not change throughout the main stage. In particular, for the homogeneous case we follow the epsilon-greedy paradigm from multi-armed bandits, choosing a random arm with fixed probability $\epsilon$ and an estimated best arm otherwise. [^10] The exploration probability $\epsilon$ is chosen in a non-standard way depending on the prior, and the estimated best arm is determined by the warm-up data. Third, counterfactual estimates are formed using the standard machinery of Inverse Propensity Scores (IPS). In contrast, our negative results apply to a wide class of mechanisms, allowing arbitrary dependence on the past data (in most rounds), arbitrary changes throughout the main stage, and arbitrary counterfactual estimates. The substance of our analysis is connecting both upper and lower bounds to the single-round benchmark.
We emphasize that our contribution spans the whole “journey”: from the new model to the benchmark to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. Conceptually, we connect the dynamic process and statistical estimators in the mechanism with a static information-design problem in the benchmark. We focus on the asymptotically-optimal analysis of the main stage, given with some “constant” (prior-dependent) amount of warm-up data. While we characterize the sufficient amount of warm-up data, we suspect it could be improved.
@@
Map of the paper. We first present and analyze our model for the case of homogeneous agents (Sections 3 and 4). This is to introduce the core model and ideas in a more accessible way, disentangling them from the considerable extra complexity (ideas, technicalities, and heavier notation) needed for the general case. The main case of heterogeneous agents is defined and treated in Section 5. The extension with estimated type frequencies is in Section 6. The technical machinery in Sections 4, 5 and 6 is introduced gradually and heavily reused from one section to the next.
@@ 2 Related work
Economic design. Using information asymmetry and commitment power to incentivize agents to take a particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor. The key distinction is that the designer’s utility is (typically) directly derived from each agent’s actions. In contrast, our statistical objective depends on the sequence of sampling distributions and does not add up across agents.
@@
Difficulties in patient recruitment are widely acknowledged as a significant barrier (e.g.,  Adams et al.,, 1997; Grill and Karlawish,, 2010). Its reasons are well-studied, e.g., see surveys (Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). Patients’ preferences among treatments are listed among the major factors, along with other factors such as insufficient trust, logistical difficulties, or aversion to randomization and loss of control. Prior work suggested increasing the probability of the “new” treatment as a possible remedy (Vozdolska et al.,, 2009; Jenkins and Fallowfield,, 2000; Pocock,, 1979; Karlawish et al.,, 2008), albeit without a formal model of participation incentives and without theoretical guarantees. These papers focus on the scenario when the new treatment appears ex-ante preferable compared to non-participation, and do not revise the treatment probabilities as the new data arrives. Other suggested remedies, focusing on eligibility criteria, patient outreach, trial logistics, etc., are not as relevant to this paper.
Data-adaptive designs for RCTs are not uncommon, e.g., see Chow and Chang, (2008) for a research survey, and Detry et al., (2012); Pallmann et al., (2018) for practitioners’ guidance. However, while prior work used data-adaptivity to improve efficiency (the tradeoff between the trial’s size and statistical properties), we use it to create participation incentives. Accordingly, the novelty in our RCT design is in connecting it to the analysis of incentives (and the corresponding lower bounds on the statistical objective), not in the RCT design per se. As far as the design itself, our RCT for the homogeneous-agent case follows a very standard “epsilon-greedy” approach from multi-armed bandits, whereas the design for the heterogeneous agents appears new.
Our mechanisms are superficially similar to patient preference trials (PPTs) (Torgerson and Sibbald,, 1998), in that both allow patient preferences to impact treatment allocation. The key difference is that PPTs allow the patients to directly choose the chosen treatment (potentially at the expense of internal validity), whereas in our setting a patient’s report only affects the sampling distribution (hence internal validity is not compromised). Another distinction is that we elicit patient preferences over the treatment outcomes, not over the treatments themselves.
Statistics and machine learning. Our mechanisms use a standard counterfactual estimator called inverse propensity score (IPS). While more advanced estimators are known (e.g.,  Dudík et al.,, 2014; Swaminathan and Joachims,, 2015; Wang et al.,, 2017), IPS is optimal in some worst-case sense (see Section 3.4 and Appendix D), which suffices for our purposes.
@@ 3.3 Economic model: agents, beliefs and incentives
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^16] We invoke agents’ beliefs because satisfying the IR property for every adversary (a.k.a. ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all agents. 171717 We called such beliefs time-invariant in the Introduction. These beliefs are misspecified, meaning that their support might not include the actual adversary (which need not be stochastic). See Related Work for background. All agents share the same belief, a common simplification in economic theory. Such beliefs represent the current medical knowledge relevant to the clinical trial, or the learnings from user interactions collected by the online platform (which the platform may choose to share with its users).
@@ 4.1 Benchmark
The inequality in (LABEL:eq:lm:pf:bench-single) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (4.2). This holds for any $\delta>0$. ∎

### v12

@@ Abstract
Keywords: exploration-exploitation tradeoff, incentive-compatibility, economic design, randomized controlled trials, multi-armed bandits
@@ 1 Introduction
Our scope. We initiate the study of incentivized participation in RCTs via information asymmetry rather than subsidies. Our objective is to incentivize participation while optimizing statistical performance. We frame it as a non-standard variant of exploration-exploitation tradeoff: the RCT always prefers to explore as uniformly as possible, whereas each patient/user prefers to exploit (choose an arm that seems best), and might not participate if the RCT is inconsistent with this preference. [^5] The standard tradeoff has an algorithm balance exploration vs exploitation to converge on the best arm. We henceforth refer to the potential RCT participants as agents.
@@
Adversarial outcomes and time-invariant beliefs. We aim for statistical guarantees that are as strong as those achievable by RCTs with per-agent randomization. Therefore, we allow the outcomes to be generated by an adversary (which motivates the need for per-agent randomization). In particular, our statistical guarantees are model-free once the data are collected.
@@
For participation incentives, we aim for guarantees that are standard in the relevant literature on economic design. While we allow for adversarial outcomes, we posit that the agents believe the outcome-generating process is much more well-behaved. Specifically, the agents believe that one’s outcome distribution is determined by the assigned arm (and one’s public type, if applicable), independent of the round. We call such beliefs time-invariant. We posit that the beliefs are shared by all agents and are known to the mechanism, as an appropriately defined Bayesian prior. Time-invariant beliefs are standard in modeling sequential decision-makers (e.g., in incentivized exploration, see Related Work).
@@
This dichotomy between adversarial outcomes and simpler, time-invariant beliefs of the agents is essential to our model. While the time-invariant beliefs need not be correct, they are reasonable as a model of reality and quite realistic as beliefs that patients or platform users might have, particularly so for clinical RCTs where they represent pre-existing medical knowledge. Beliefs whose support does not include the actual problem instance (a.k.a. misspecified beliefs) are not uncommon in economic theory, see Related Work.
Our model. We model “incentivized participation” as a mechanism design problem under a statistical objective and incentive constraints. An RCT (mechanism) sequentially interacts with self-interested agents like a multi-armed bandit algorithm: in each round, an agent arrives, the mechanism chooses an arm for this agent, and observes an outcome. The mechanism’s statistical objective is standard for RCTs: counterfactually estimate each arm’s performance on the whole population, even if the outcomes are generated by an adversary. [^7] In contrast, the standard objective for multi-armed bandits is to optimize cumulative payoffs. While there are several reasonable ways to formalize this objective, we choose one that is most convenient for our purposes: the worst-case mean-squared estimation error, denoted $\mathtt{ERR}$ (and formally defined in Section 3.2). If not for the incentive constraints for participation, the mechanism would simply choose arms independently and uniformly at random for each agent.
The role of $f$, the RCT’s metric, is as follows. Typically, $f(\omega)$ corresponds to the “societal value” of the outcome $\omega$. The RCT may also be interested in the empirical frequency of $\omega$, which is captured by setting $f={\bf 1}_{\left\{\,\omega\,\right\}}$, the indicator function that chooses this outcome. Our approach does not hinge upon a particular choice of $f$. In fact, our definition of $\mathtt{ERR}$ handles all metrics $f$ at once, and therefore applies to estimating societal value as well as outcome frequencies. In contrast, agents have subjective utilities over the outcomes that they wish to optimize, possibly different from the “societal value” or any other metric used by the RCT.
@@
We address participation incentives by imposing Bayesian Individual Rationality (BIR) constraints, a standard notion in economic theory. Specifically, the mechanism commits to some randomized policy that assigns an arm to each agent, depending on the history of interactions with the previous agents. The agent does not observe the history or the chosen arm (as discussed above). The BIR constraint states that each agent must prefer participation to her “outside option,” and it (only) needs to hold with respect to agents’ time-invariant beliefs. In particular, it suffices to provide participation incentives for the hypothetical world in which the agents’ beliefs are correct. We note that a stronger version of participation incentives — robust against an adversarially chosen outcome sequence — is impossible to achieve (cf. Footnote 17).
We consider three model variants. To isolate the core issues of incentivized participation, we first consider the fundamental case of homogeneous agents with the same beliefs and preferences. Then, the general case of heterogeneous agents endows each agent with a type. Then observable attributes such as medical history or user profile comprise one’s public type and determine one’s beliefs over outcomes. Moreover, one has unobservable preferences over the outcomes (private type) which need to be elicited in an incentive-compatible way. [^8] In practice, preference elicitation would happen during registration/opt-in process. For online platforms, private preferences and their elicitation make sense only if the RCT has explicit opt-in. The latter is uncommon, but may be needed if the RCT involves high-stakes interactions (such as buying a house) or is potentially controversial. The third model variant provides the mechanism with estimated type frequencies over the agent population. We address situations in which some types are inherently (much) more “difficult” to handle than others. When such “difficult” types are rare, we leverage the type frequencies to mitigate their harmful influence.
@@
Our results. We focus on mechanisms with simple two-stage designs: a relatively short warm-up stage of predetermined duration, followed by the main stage that needs to be counterfactually estimated. The data from the warm-up stage (warm-up data) creates information asymmetry in the main stage which creates participation incentives, whereas the main stage determines the counterfactual estimates.  [^9] The main stage may, in principle, also contribute to participation incentives. We found this capability unnecessary for our optimal mechanisms, but our lower bounds allow it. RCTs with a warm-up stage are very common in practice: e.g., phase-I clinical trials to calibrate the dosage and “pilot experiments” in online platforms.
@@
We obtain a near-optimal solution for each model variant described above, from homogeneous agents to heterogeneous agents to estimated type frequencies. First, we put forward an idealized benchmark which focuses on a single round of our problem. Then, we design a mechanism whose estimation error $\mathtt{ERR}$ (in our actual, multi-round model) is upper-bounded in terms of this single-round benchmark. Finally, we prove that no mechanism achieves better estimation error, up to constant factors. [^10] The upper and lower bounds show that $\mathtt{ERR}$ is on the order of $B/T$, where $B$ is the benchmark. This is a different notion of optimality compared to bandits (where one usually optimizes regret, i.e., the $o(T)$ additive term). While the benchmark, the mechanism, and the lower bound are different for each model variant, many ideas carry over from one model variant to the next.
@@
Our mechanisms have a simple structure. First, their type-dependent “sampling distributions” (from which the arms are sampled) do not adapt to the data collected in the main stage and do not change throughout the main stage. Second, counterfactual estimates are formed using the standard machinery of Inverse Propensity Scores (IPS). In contrast, our negative results apply to a wide class of mechanisms, allowing arbitrary dependence on the past data (in most rounds), arbitrary changes throughout the main stage, and arbitrary counterfactual estimates. Thus, our “simple” mechanisms are proved near-optimal among arbitrary “complex” mechanisms.
Significance. Our contribution spans the whole “journey”: from the new model to the benchmark to the mechanisms to the matching upper/lower bounds, and from homogeneous agents to heterogeneous agents to estimated type frequencies. In particular, the inherent complexity of each problem variant is expressed in terms of the respective single-round benchmark, a much simpler, “static” object amenable to further analysis; modeling/definitions are absolutely crucial for that. The substance of our analysis is connecting the dynamic process and statistical estimators in the mechanism with a static information-design problem in the benchmark. We emphasize homogeneous agents to introduce the core model and ideas in a more accessible way. The general case requires a heavier setup, non-trivial mechanisms (within the simple structure outlined above), and a much more technical analysis.
The non-adaptivity of our mechanisms is desirable in applications. Indeed, clinical RCTs in practice typically do not adapt to data, except possibly at a few pre-determined “review points”, [^11] This is because treating the patients and/or observing their outcomes may take a long time, and the doctors may find it difficult to coordinate the treatments across the patients. and neither do most RCTs in online platforms due to engineering considerations. In fact, many/most patients (resp., platform users) may need to be treated in parallel, and this is feasible for the main stage in our mechanisms. On the other hand, data-adaptive data collection allowed by our lower bounds is not out of scope, either, both for online platforms (e.g.,  Agarwal et al.,, 2016, 2017) and for clinical RCTs (see Related Work). We conjecture it may help in addressing some of the open questions outlined in Conclusions.
We focus on the asymptotically-optimal analysis of the main stage, illuminating the inherent power and limitations of information-driven participation incentives. Our guarantees for the warm-up stage make a point (that a constant-sized warm-up suffices for any $T$) and emphasize interpretability, but we suspect they can be improved.
@@
Map of the paper. We first present and analyze our model for the case of homogeneous agents (Sections 3 and 4). The main case of heterogeneous agents is defined and treated in Section 5. The extension with estimated type frequencies is in Section 6. The technical machinery in Sections 4, 5 and 6 is introduced gradually and heavily reused from one section to the next.
@@ 2 Related work
Economic design. Narita, (2021) incentivizes participation in clinical RCTs when the patients are certain about the arms’ payoffs and have a fixed outside option. Accordingly, information asymmetry (and learning over time to create it) plays no role in participation incentives. Certain beliefs are a degenerate case from our perspective (which allows and leverages distributional beliefs).
Using information asymmetry and commitment power to incentivize agents to take a particular action is a common goal in information design (Bergemann and Morris,, 2019; Kamenica,, 2019). However, this literature tends to have a very different technical flavor. The key distinction is that the designer’s utility is directly derived from each agent’s actions. In contrast, our statistical objective depends on the sequence of sampling distributions and does not add up across agents.
@@
Chassang et al., (2012) apply economic design to RCTs for a different purpose, focusing on agents’ ability to strategically choose their effort level. They use principal-agent framework to disentangle the effects of treatment and effort without compromising external validity. Within this framework, incentives are created via monetary contracts.
@@
Difficulties in patient recruitment are widely acknowledged as a significant barrier (e.g.,  Adams et al.,, 1997; Grill and Karlawish,, 2010). The reasons are well-studied, e.g., see surveys (Ross et al.,, 1999; Mills et al.,, 2006; Rodríguez-Torres et al.,, 2021). Patients’ preferences among treatments are listed among the major factors, along with other factors such as insufficient trust, logistical difficulties, aversion to randomization, and loss of control. Prior work suggested increasing the probability of the “new” treatment as a possible remedy (Vozdolska et al.,, 2009; Jenkins and Fallowfield,, 2000; Pocock,, 1979; Karlawish et al.,, 2008), albeit without a formal model of participation incentives and without theoretical guarantees. These papers focus on the scenario when the new treatment appears ex-ante preferable compared to non-participation, and do not revise the treatment probabilities as new data arrive. Other suggested remedies focusing on eligibility criteria, patient outreach, trial logistics, etc., are not as relevant to this paper.
Data-adaptive designs for RCTs are not uncommon and have a long history in biostatistics, e.g., see Chow and Chang, (2008) for a research survey, and Detry et al., (2012); Pallmann et al., (2018) for practitioners’ guidance. While prior work used data-adaptivity to improve efficiency (the tradeoff between the trial’s size and statistical properties), we use it to create participation incentives. Accordingly, the novelty in our RCT design is in connecting it to the analysis of incentives (and the corresponding lower bounds on the statistical objective), not in the RCT design per se. That said, our RCT for the homogeneous-agent case is based on a standard “epsilon-greedy” approach from multi-armed bandits, whereas our design for heterogeneous agents appears new.
Our mechanisms are superficially similar to patient preference trials (PPTs) (Torgerson and Sibbald,, 1998), in that both allow patient preferences to impact treatment allocation. The key difference is that PPTs allow the patients to directly choose the treatment (potentially at the expense of internal validity), whereas in our setting a patient’s report only affects the sampling distribution (hence internal validity is not compromised). Another distinction is that we elicit patient preferences over the treatment outcomes, not over the treatments themselves.
Statistics and machine learning. Our mechanisms use a standard counterfactual estimator called inverse propensity score (IPS). While more advanced estimators are known (e.g.,  Dudík et al.,, 2014; Swaminathan and Joachims,, 2015; Wang et al.,, 2017), IPS is optimal in some particular worst-case sense (see Section 3.4 and Appendix D), which suffices for our purposes.
@@ 3.3 Economic model: agents, beliefs and incentives
The mechanism must incentivize participation, i.e., satisfy Individual Rationality (IR). We require Bayesian IR in expectation over the agents’ beliefs over the adversaries. [^17] We invoke agents’ beliefs because satisfying the IR property for every adversary (a.k.a. ex-post IR) is clearly impossible, except for a trivial mechanism that always chooses the agent’s outside option. Crucially, we posit that agents’ beliefs are over stochastic adversaries, i.e., the agents believe that a given treatment yields the same outcome distribution in all agents.  [^18] We called such beliefs time-invariant in the Introduction. These beliefs are misspecified, meaning that their support might not include the actual adversary (which need not be stochastic). See Related Work for background. All agents share the same belief, a common simplification in economic theory. Such beliefs represent the current medical knowledge relevant to the clinical trial, or the learnings from user interactions collected by the online platform (which the platform may choose to share with its users).
@@ 4.1 Benchmark
$U_{\mathcal{P}}(\sigma^{\prime})$ $=\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,(1-\epsilon+\epsilon/n)\cdot U_{\psi}(a^{*}(\psi))+\frac{\epsilon}{n}\cdot{\sum_{a\neq a^{*}(\psi)}}U_{\psi}(a)\,\right]$
$\geq\operatornamewithlimits{\mathbb{E}}_{\psi\sim\mathcal{P}}\left[\,\sigma_{a^{*}(\psi)}(\psi)\cdot U_{\psi}(a^{*}(\psi))+\sigma_{a}(\psi)\cdot{\sum_{a\neq a^{*}(\psi)}}U_{\psi}(a)\,\right]$ (4.3)
$=U_{\mathcal{P}}(\sigma)\geq U_{\mathtt{out}}.$
The inequality in (4.3) holds since $\sigma_{a}(\psi)\geq\epsilon/n$, so for each state the right-hand side shifts probabilities towards arms with weakly smaller expected utility. It follows that policy $\sigma^{\prime}$ is BIR. So, this policy gets within $\delta$ from the supremum in (4.2). This holds for any $\delta>0$. ∎